# Convergence Rate for the Last Iterate of Stochastic Gradient Descent Schemes

### v3

@@ 1. Introduction
Algorithms of the form (1.2) have been studied extensively as part of a field called stochastic approximation ([10], [13]). However, the almost sure convergence rate for the last iterate of SGD is studied only recently. The most recent work in the parametric setting are done by Weissmann et al in [15]. Weissmann et al. considers SGD and SHB where properties of the objective function $F$ is only known locally, i.e. properties apply in some neighborhood of a point [15, assumption 2.1.(ii)]. Under the local gradient domination [15, definition 2.2] $\|\nabla F(w_{t})\|^{2}\geq F(w_{t})-F_{*}$, Weissmann et al proved the existence of neighborhoods $\pazocal{U}_{1},\pazocal{U}$ of critical points in the parameter space $\pazocal{W}$ for which if $w_{1}\in\pazocal{U}_{1}$, then the event $\Omega{U}=\{\pazocal{U}\,\text{is absorbing for}\,w_{t}\}$ is highly probable and $(F(w_{t})-F_{*})\mathbbm{1}_{\Omega{U}}=o(t^{-1+\epsilon})$ [15, table 1, theorem 5.1]. The set $\pazocal{U}$ is defined as a neighborhood of the set of minimizers with a fixed radius. We remark that Weissmann et al also examines the case when global properties of $F$ are known [15, table 1].
When global properties of the objective function are known, there are other recent works as well. Liu and Yuan proved almost sure convergence rates for SGD, SHB, and SNAG (Stochastic Nesterov’s Accelerated Gradient) by assuming that the objective function is either globally non-convex, convex, or strongly convex [9]. Liu and Yuan produced the rate $F(w_{t})-F_{*}=o(t^{-1+\epsilon})$ for strongly convex $F$ and $\min_{1\leq i<t-1}\|\nabla F(w_{i})\|^{2}=o(\sum_{i=1}^{t-1}\alpha_{i})^{-1}$ for non-convex $F$ [9, theorem 6, theorem 8]. The strongly convex rate is optimal according to Agarwal et al [1]. For a convex objective function $F$, Liu and Yuan proved a rate of $F(w_{t})-F_{*}=o(t^{-\frac{1}{3}+\epsilon})$, which is not as close to the optimal rate of $t^{-1/2}$ in expectation [9, remark 14]. Liu and Yuan used the Robbins-Siegmund theorem ([13, theorem 1], [9, proposition 2]) to prove their convergence rate results. We remark that both SGD and SHB produce the same rate according to Liu and Yuan, i.e. the momentum parameter $\beta$ does not affect the convergence rate [9, theorem 8, theorem 13].
Intuitively, the momentum parameter $\beta$ in (1.1) serves to counter slowdown when the gradient is near zero. As such, it makes sense to adjust $\beta$ depending on the gradient and step sizes. Sebbouh et al studied such a case in [14]. In particular, they proved that the last iterate of SHB-IMA (Iterate Moving Average) can achieve $o(t^{-1+\epsilon})$ rate in the convex setting when $\beta$ is allowed to vary as a function of $t$ in the ‘overparametrized’ case [14, table 1, corollary 14, definition 3]. On the other hand, in the usual parametrized case, given step size of order $O(t^{-\frac{1}{2}-\epsilon})$, Sebbouh et al achieved $o(t^{-\frac{1}{2}+\epsilon})$ rate [14, corollary 17]. Similar to Liu and Yuan’s work [9], Sebbouh et al used the Robbins-Siegmund theorem [14, lemma 6] to prove these rates.
In another work [8], Lei, Shi, and Guo obtained the last iterate almost sure convergence rate of $o(t^{\max(p-1,1-p(\gamma+1))+\epsilon})$ given step size $O(t^{-p})$ with $p\in(\frac{1}{1+\gamma},1)$ for a convex objective function, whose gradient is $\gamma$-Hölder with SGD in the non-parametric setting [8, theorem 6]. Moreover, in [8, corollary 11], Lei, Shi, and Guo proved that SGD converges at rate $F(w_{t+1})-F_{*}=O(t^{\max(p-1,1-p(\gamma+1))}\,\log^{2}(t/\delta))$ with probability at least $1-\delta$ if $p\in(\frac{1}{1+\gamma},1)$. In particular, when $p=2/3$ and $\gamma=1$, their work matches Liu and Yuan’s with $O(t^{-\frac{1}{3}+\epsilon})$ [8, theorem 6, corollary 11]. Interestingly, Lei, Shi, and Guo did not use Robbins-Siegmund theorem. However, their approach is similar to that of Robbins and Siegmund, where a supermartingale is constructed and the convergence of the iterates is implied by that of the supermartingale [8, section 4.2].
Another recent work regarding almost sure convergence rate for SGD without employing the Robbins-Siegmund theorem, is done by Nguyen et al [11]. The authors used a weaker, super-martingale convergence theorem [11, lemma 5] to prove $w_{t}\rightarrow w_{*}$ a.s.-$\mathbb{P}$ and $\mathbb{E}\|w_{t}-w_{*}\|^{2}=o(t^{-1})$ in the strong convex case [11, thorem 1,2]. Such a rate is indeed optimal according to [1] since if $F$ is convex with $\nabla F$ Lipschitz, by lemma C.4, we get $F(w_{t})-F_{*}\lesssim\|w_{t}-w_{*}\|^{2}=o(t^{-1})$.
@@ 1.1. Contribution
As remarked in Liu and Yuan’s paper [9, remark 14], the rate $O(t^{-\frac{1}{3}+\epsilon})$ for the convex case is not close to the optimal rate in expectation $O(t^{-\frac{1}{2}})$ proved by Agarwal et al in [1]. Motivated by this remark, we study the almost sure convergence for the last iterate of SGD and SHB where the objective function $F$ is globally convex and non-convex with $\nabla F$ being $(\gamma,L)$-Hölder. With these conditions, our contribution is as follows:
  1. 1.
    Instead of relying on Robbins-Siegmund, nor any stochastic approximation theory, we use only Gronwall’s inequality without using any martingale convergence theorem to prove that $\min_{0\leq s\leq t}\|\nabla F(w_{t})\|^{2}=o(t^{p-1})$ and $F(w_{t})-F_{*}=o(t^{(2\gamma/(1+\gamma))\max(p-1,1-(1+\gamma)p)-\epsilon})$, as well as $\min_{s\leq t}F(w_{s})-F_{*}=o(t^{p-1})$ for non-convex and convex objectives respectively given step size of $\alpha_{t}=O(t^{-p})$ with $p\in(\frac{1}{1+\gamma},1)$ using similar assumptions as in [9], except that $\nabla F$ is assumed to be $\gamma$-Hölder instead of Lipschitz. By using non-martingale convergence theory, our argument is less technical. We also present insights from our study in section 2.2 and 2.3.
  2. 2.
    We proved that with probability at least $1-\delta$, SHB (1.1) with a fixed momentum parameter $\beta\in(0,1)$ converges at rate $F(w_{t})-F_{*}=O(t^{\max(p-1,-2p+1)}\log^{2}\frac{t}{\delta})$ when the gradient of the objective function $\nabla F$ is Lipschitz and $F$ is convex. Our result is similar to [8] and is consistent with our own findings (see table 1). While our result is restricted to the case when $\beta$ is fixed, such a result provides baseline SHB performance which can be compared to other algorithms, e.g. SHB-IMA (Iterated Moving Average). There may be arguments to use baseline (simpler) SHB variant to save computation time. To the author’s knowledge, this is the first high probability convergence rate result for the SHB.
  3. 3.
    We use a slightly weaker assumption for $\nabla F$ compared to Liu and Yuan [9, assumption 4] that it is only $\gamma$-Hölder with $\gamma\in(0,1]$, following Lei, Shi, and Guo [8, assumption 1]. In this manner, we provide a unified analysis for ‘smooth’ objective functions.
@@ 1.2. Acknowledgment
Algo Cost Function w/ $(\gamma,L)$-smoothness Step Size Decay Rate $p$ in $O(t^{-p})$ Convergence Rate for $F(w_{t})-F_{*}$ (convex case) or $\|\nabla F(w_{t})\|^{2}$ (non-convex) Statement
SGD Convex, $\gamma\in(0,1]$ $(\frac{1}{1+\gamma},1)$ $o(t^{\max(p-1,1-(\gamma+1)p)-\epsilon})$ theorem 2.7, [8, theorem 6].
$(\frac{1}{2},1)$ $O(t^{\max(p-1,1-(\gamma+1)p)}\log^{2}\frac{t}{\delta})$ with prob. $1-\delta$ [8, corollary 11].
Convex, $\gamma=1$ $(\frac{2}{3},1)$ $o(t^{p-1})$ [9, theorem 13].
Non-convex, $\gamma\in(0,1]$ $(\frac{1}{1+\gamma},1)$ $\min_{0\leq s\leq t}\|\nabla F(w_{t})\|^{2}=o(t^{p-1})$ corollary 3.3.
Non-convex, $\gamma=1$ $(\frac{1}{2},1)$ $\min_{1\leq i\leq t-1}\|\nabla F(w_{t})\|^{2}=o\left(\frac{1}{\sum_{i=1}^{t-1}\alpha_{i}}\right)$ [9, theorem 6].
SHB Convex, $\gamma\in(0,1]$ $(\frac{1}{1+\gamma},1)$ $o(t^{\frac{2\gamma}{1+\gamma}\max(p-1,1-(\gamma+1)p)-\epsilon})$ theorem 2.7.
Convex, $\gamma=1$ $(\frac{2}{3},1)$ $o(t^{p-1})$ [9, theorem 13].
$(\frac{1}{2},1)$ $O(t^{\max(p-1,1-2p)}\log^{2}\frac{t}{\delta})$ w/ prob. $1-\delta$ theorem 2.10.
Non-convex, $\gamma\in(0,1]$ $(\frac{1}{1+\gamma},1)$ $\min_{1\leq s\leq t}\|\nabla F(w_{t})\|^{2}=o(t^{p-1})$ corollary 3.6.
Non-convex, $\gamma=1$ $(\frac{1}{2},1)$ $\min_{1\leq i\leq t-1}\|\nabla F(w_{t})\|^{2}=o\left(\frac{1}{\sum_{i=1}^{t-1}\alpha_{i}}\right)$ [9, theorem 8].
SHB-IMA Convex, $\gamma=1$ $\frac{1}{2}-\epsilon$ $o(t^{-\frac{1}{2}+\epsilon})$ [14, corollary 17].
Table 1. Summary of Results.
@@ 2.1. Notation and Assumptions
$E_{n}$ eventually The event $\bigcup_{m=1}^{\infty}\bigcap_{n\geq m}E_{n}$ ($\exists m>0$ such that $E_{n}$ occurs for all $n\geq m$).
@@
Assumption 2.3 is also used in [8]. Such an assumption can be satisfied when $\pazocal{Z}$ is bounded. See discussions in [8]. Note that by lemma C.6.3b, if $\ell(z,\cdot)$ is convex and assumption 2.1 holds, then the ‘noise’ $\|\delta m_{t}\|^{2}$ defined in (1.2) is bounded by a scalar multiple of $\|w_{t}-w_{*}\|^{2\gamma}+1$. This assumption will only be used to produce probability estimates for the convergence rate of SHB iterates.
2.2. Main Result 1: Almost Sure Convergence Rate
With assumption 2.1 and 2.2, in estimating $F(w_{t})-F_{*}$, one naturally arrives at a recursive inequality. In this regard, our key argument is a deterministic, weaker version of the Robbins-Siegmund theorem [13] and a lemma which deduce the rate.
@@ 2.2. Main Result 1: Almost Sure Convergence Rate
We include the ‘best case’ scenario (prefixed with $\min_{0\leq s\leq t}$) as part of our result. Best case scenario may be of less interest, but still provides information on how the rate compares with the ‘normal case’. In particular, it shows that the ‘optimal’ $t^{-1/2}$ rate for $\gamma=1$ is achievable if $p=\frac{1}{2}+\epsilon$. Further, in the application, one may stop the algorithm using a stopping time, in which case one may be interested in the ‘best case’ scenario.
Remark 2.9.
The slowdown factor $r_{\gamma}$ due to the smoothness parameter $\gamma$ appears only with momentum $\beta>0$ and $\gamma<1$. When $\gamma<1$, intuitively, the algorithm is slowed down as can be seen in the last term of (2.1). The fact that the slowdown factor $r_{\gamma}$ is uniform for all $\beta>0$ is interesting as empirical evidence showed that larger $\beta$ contributes to more slowdown [3]. There are also other works on non-acceleration of heavy ball [5]. We do not quite understand at the moment of these phenomena and how it ties to the slowdown factor $r_{\gamma}$. In addition, it is counter-intuitive since momentum is supposed to make convergence faster when the gradient is small. Examining (3.13), when there is momentum, the gradient $\|\nabla F(w_{t}+\frac{\beta}{1-\beta}(w_{t}-w_{t-1}))\|^{2}$ appears as an additive term in the upper estimate for $F(w_{t})-F_{*}$ and lemma C.4 suggests that for $\gamma<1$, this slows down the rate.
2.3. Main Result 2: Convergence Rate with High Probability
A key requirement to obtain fast asymptotic convergence is that $F(w_{t})-F_{*}$ is decreasing eventually. In other words, there exists $T>0$ such that $F(w_{t})$ is decreasing for all $t>T$. Such is the case in deterministic GD [4, lemma 2.28, eq. (14), algorithm 3.2]. In fact, smoothness of $F$ is sufficient (without convexity). Intuitively, in the stochastic GD, this would hold when $F$ is convex if the step size decay rate is controlled accordingly. However, even after taking expectation, this is difficult to achieve. Indeed, applying the expectation to (3.2) with [4, lemma 2.29] and proposition 3.1 yields an inequality of the form
$\mathbb{E}[F(w_{t+1})]-\mathbb{E}[F(w_{t})]$ $\leq-\alpha_{t}\bigg(\mathbb{E}\|\nabla F(w_{t})\|^{2}-\alpha_{t}^{\gamma}\left(A\,\mathbb{E}[F(w_{t})-F_{*}]+B\,\|\nabla F(w_{t})\|^{2}+C\right)\bigg)$
$\leq-\alpha_{t}\bigg(\min_{0\leq s\leq t}\mathbb{E}\|\nabla F(w_{s})\|^{2}-\alpha_{t}^{\gamma}\left(C+(A+2BL)\,\sup_{t\geq 0}\mathbb{E}[F(w_{t})-F_{*}]\right)\bigg).$
From [8, theorem 6], convex $F$ gives $\mathbb{E}\|\nabla F(w_{t})\|^{2}\leq\mathbb{E}[F(w_{t})-F_{*}]=O(t^{p-1})$ for $p>2/(2+\gamma)$. Since $p>1/(1+\gamma)$, we have $t^{p-1}>t^{p\gamma}$ which suggests that on average, $F(w_{t})$ decreases eventually.
We address this issue in our second result on SHB. Instead of looking at one-step iterates, we look at $F(w_{t})$ over a time period $t\in[\tau,T+1]$ in (4.35) by expressing
$F(w_{T+1})$ $\leq F(w_{\tau})+\sum_{t=\tau}^{T}\left(\beta\,\langle\nabla F(w_{t}),\,w_{t}-w_{t-1}\rangle+\frac{L}{1+\gamma}\,\|w_{t+1}-w_{t}\|^{1+\gamma}\right)$
$-\sum_{t=\tau}^{T}\left(\alpha_{t}\,\|\nabla F(w_{t})\|^{2}-\alpha_{t}\,\langle\nabla F(w_{t}),\,\delta m_{t}\rangle\right).$
The first summation on the right hand side decays quickly. The key is to control the last summation. Following the same idea as Lei, Shi, Guo in [8], we use concentration inequality of Bernstein and Azuma-Hoeffding to estimate the last term on the right hand side involving the martingale difference $\alpha_{t}\,\langle\nabla F(w_{t}),\,\delta m_{t}\rangle$. Effectively, the martingale difference term can be made smaller than $\alpha_{t}\,\|\nabla F(w_{t})\|^{2}$ so that the last term is effectively negative (more precisely, it is replaced with a quickly decaying term. See lemma 4.10). This suggests that with high probability, the ‘decreasing’ effect on $F(w_{t})$ (the effectiveness of SGD/SHB) is shown over a larger time step, which is a slowdown compared to determinsitic GD [4, theorem 3.4, lemma 2.28].
From the previous paragraph, applying martingale concentration inequalities on $\langle\nabla F(w_{t}),\,\delta m_{t}\rangle$ requires controlling the noise $\delta m_{t}$. By definition, $\delta m_{t}$ is comprised of $\nabla F(w_{t})$ and $\nabla\ell(Z_{t},w_{t})$, which by lemma C.4, is controlled by $\|w_{t}-w_{*}\|$. Following [8], in proposition 4.5, we show that up to time $T>0$, $\|w_{t}-w_{*}\|$ is at most $\log(T/\delta)$ with probability $1-\delta$. Since $\alpha_{t}$ decays with order $O(t^{-p})$, the martingale difference $\alpha_{t}\,\langle\nabla F(w_{t}),\,\delta m_{t}\rangle$ is of the order $\alpha_{t}$ up to a logarithmic factor.

### v4

@@ 1. Introduction
SHB (1.1) has been studied quite extensively, especially when $F$ is coercive, convex, and smooth with bounded second derivative and with finite variance in the noise. Specifically, under such conditions, results on almost sure convergence, scaling limit in distribution, and $L^{2}$ convergence rates have been derived [3, theorem 1-2,4,6]. Similar results can also be found for the special case (1.2). Such a scheme is a class of Robbins-Monro procedures [14]. See also [6]. Under some conditions on the noise and $\nabla F$, Robbins-Monro type algorithms converge almost surely [10, section I.1, thm I.9]. Such a statement relies on the Robbins-Siegmund theorem [13, theorem 1], which is a convergence theorem for non-negative almost super-martingale. We remark that other methods to treat the almost sure convergence of Robbins-Monro algorithm exists, e.g. the Ordinary Differential Equation (ODE) method, which relies on the stability theory of ODE. See [10, section I, page 11].
Going from classical theory, there are two important directions to pursue. The first belongs to the case where $F$ is non-convex with Lipschitz gradient (weakening the convexity and smoothness assumptions on $F$) under weakened assumption on the noise. The second belongs to almost sure convergence rates. Despite many classical results, strong convergence behavior of stochastic gradient schemes under general convexity assumptions on $F$ (e.g. non-convex with saddle points) under weaker noise assumption has been studied fairly recently. Liu and Yuan provided sufficient conditions on $F$ without assuming global uniform variance bound on the estimator $\|\nabla\ell\|$ for SHB and SNAG (Stochastic Nesterov Accelerated Gradient) to avoid saddle points almost surely [9]. In addition, the authors of [9] expanded the application of Robbins-Siegmund theorem for proving almost sure convergence rates for SGD, SHB, and SNAG for non-convex, convex, and strongly convex objectives whose gradient is Lipschitz [9, table 1]. We remark that other recent convergence rate results include [1, 8, 11, 12, 15, 16].
When $F$ is convex, more results are available. The optimal convergence rate $O(t^{-1/2})$ for stochastic first order oracle was proved for bounded, convex, and Lipschitz cost functions in [1, thm 1-2]. A constructive proof for such an optimal rate using SHB is done in [15]. More precisely, Sebbouh et al. showed that there exists $\{(\alpha_{t},\beta_{t}):t\geq 0\}$ such that eq. 1.1 attains the optimal convergence $O(t^{-\frac{1}{2}+\epsilon})$ for convex cost functions with Lipschitz gradient [15, theorem 13]. In this case, the momentum $\beta_{t}$ is a function of the step size $\alpha_{t}$. When the step-size is chosen to be $\alpha_{t}=O(t^{-1/2})$, the momentum takes values $\beta_{t}=O(0.5t/(t+1))$. Intuitively, when the minimizer is still far away in the early iterations, the algorithm is approximately an SGD, but gradually transitions to SHB with large momentum as $t\rightarrow\infty$. This shows that the choice of $\beta_{t}$ is crucial since the almost sure asymptotic performance of SHB is the same as SGD when $\beta_{t}=\beta$ is held constant [9]. In the non-parametric setting, Lei, Shi, and Guo provided almost sure convergence rate and fluctuation for SGD ($\beta_{t}=0$) when $\nabla F$ is only $\gamma$-Hölder [8]. More precisely, SGD achieves an almost sure convergence rate $o(t^{\max(p-1,1-p(\gamma+1))+\epsilon})$ given step size $O(t^{-p})$ with $p\in(\frac{1}{1+\gamma},1)$ for a convex objective whose gradient is $\gamma$-Hölder [8, theorem 6]. Moreover, SGD converges at rate $F(w_{t+1})-F_{*}=O(t^{\max(p-1,1-p(\gamma+1))}\,\log^{2}(t/\delta))$ with probability at least $1-\delta$ if $p\in(\frac{1}{1+\gamma},1)$ [8, corollary 11].
@@ 1.1. Contribution
We will restrict the scope of our study to the case where $\beta_{t}:=\beta$ is constant in (1.1) and $\ell(z,\cdot)$ is convex and $\nabla\ell(z,\cdot)\in\pazocal{C}^{\gamma}(\mathbb{R}^{d})$ with Hölder constant $L>0$ and $\gamma\in(0,1]$. Further, we assume that there exists a global minimum $F_{*}:=\min_{w\in\mathbb{R}^{d}}F(w)\geq 0$ and a global minimizer $w_{*}\in\mathbb{R}^{d}$. In the landscape of convergence rate theory for stochastic gradient schemes, the survey of prior works provided above indicates that the case of a convex objective whose gradient is $\gamma$-Hölder for SHB is unexplored. This is where our contribution lies.
There are three main contribution from our work. First, we provide an alternative method to prove convergence rate using Gronwall’s inequality and Doob’s martingale convergence theorem instead of using Robbins-Siegmund theorem (cf. proposition 2.4 and lemma 2.5). Second, we provide an almost sure convergence rate result for SHB for convex objective whose gradient is $\gamma$-Hölder. As stated previously, this case is unexplored. Roughly stated, the result says that given a convex objective with $\gamma$-Hölder gradient, SHB with momentum $\beta\in(0,1)$ and learning rate (step size) $t^{-p}$ with $p\in(\frac{1}{1+\gamma},1)$ converges to a global minimum almost surely with rate $\frac{2\gamma}{1+\gamma}\max(p-1,1-p(\gamma+1))$ (cf. theorem 2.7). Finally, we also provide convergence rate in probability for the case $\gamma=1$ (cf. theorem 2.9). This is also unexplored as the case $\gamma=1$ was only done for SGD. Our result is consistent with [9, 8] for the case $\gamma=1$. Table 1 compares our convergence rate results with prior work.
Algorithm Cost Function and Learning Rate Convergence Rate
SGD-np Convex, $\gamma$-Hölder gradient $O(t^{\max(p-1,1-p(\gamma+1))})$
[8] $O(t^{-p}),p\in(\frac{1}{1+\gamma},1)$ a.s. and h.p..
SHB Convex, Lipschitz gradient
[15] $O(t^{-\frac{1}{2}})$, $\beta_{t}\in O(\frac{t}{1+t})$ $O(t^{-\frac{1}{2}+\epsilon})$ a.s..
SGD, SHB, SNAG Convex, Lipschitz gradient
[9] $O(t^{-\frac{2}{3}})$ $O(t^{-\frac{1}{3}})$ a.s..
SHB, $\beta\in(0,1)$ Convex, $\gamma$-Hölder gradient $O(t^{\frac{2\gamma}{1+\gamma}\max(p-1,1-p(\gamma+1))})$ a.s.
$O(t^{-p}),p\in(\frac{1}{1+\gamma},1)$ $O(t^{\max(p-1,-2p+1)})$ h.p., $\gamma=1$.
Table 1. List of Prior works on Convergence Rates. The notation $f(t)=O(g(t))$ is defined as the existence of a positive constant $c>0$ such that $f(t)\leq c\,g(t)$ for sufficiently large $t>0$. Such definition along with other measure of convergence rates such as $f(t)=\Theta(g(t))$ and $f(t)=o(g(t))$ can be found in table 2.
@@ 2.1. Notation and Assumptions
$E$ a.s. The event $E$ occurs almost surely, i.e. $\mathbb{P}(E)=1$.
$E$ h.p. The event $E$ occurs with high probability, i.e. $\exists\delta>0,~\mathbb{P}(E)\geq 1-\delta$.
@@
Assumption 2.3 is also used in [8]. Such an assumption can be satisfied when $\pazocal{Z}$ is bounded. See discussions in [8]. Note that by lemma C.6.3b, if $\ell(z,\cdot)$ is convex and assumption 2.1 holds, then the ‘noise’ $\|\delta m_{t}\|^{2}$ defined in (1.2) is bounded by a scalar multiple of $\|w_{t}-w_{*}\|^{2\gamma}+1$. This assumption is used to produce probability estimates using concentration inequalities for the convergence rate of SHB iterates.
2.2. Almost Sure Convergence Rate
Using assumptions 2.1 and 2.2, one may estimate the iterates $F(w_{t})-F_{*}$. Indeed, under $\gamma$-Hölder smoothness assumption, we have $F(y)-F(x)\leq\langle\nabla F(x),\,y-x\rangle+(L/(1+\gamma))\,\|y-x\|^{1+\gamma}$ for all $x,y\in\mathbb{R}^{d}$. If one substitutes $y=w_{t+1}$ and $x=w_{t}$, then for the SGD case, we get an estimate for the iterates of $F(w_{t})$ almost surely:
$F(w_{t+1})-F(w_{t})\leq-\alpha_{t}\,\|\nabla F(w_{t})\|^{2}+\alpha_{t}\,\langle\nabla F(w_{t}),\,\delta m_{t}\rangle+\frac{L}{1+\gamma}\,\alpha_{t}^{1+\gamma}\|\nabla\ell(Z_{t},w_{t})\|^{1+\gamma}.$ (2.5)
A natural approach to deal with (2.5) is to apply $\mathbb{E}[\cdot\,|\,\mathscr{F}_{t}]$. Upon such application, the martingale difference term vanishes and the negative term $\alpha_{t}\|\nabla F(w_{t})\|^{2}$ can be dropped. So the iterates of $\mathbb{E}[F(w_{t+1})-F_{*}\,|\,\mathscr{F}_{t}]$ looks like
$\mathbb{E}[F(w_{t+1})-F_{*}\,|\,\mathscr{F}_{t}]\leq[F(w_{t})-F_{*}]+\frac{L}{1+\gamma}\,\alpha_{t}^{1+\gamma}\,\mathbb{E}[\|\nabla\ell(Z_{t},w_{t})\|^{1+\gamma}\,|\,\mathscr{F}_{t}].$ (2.6)
[…]
@@ 2.2. Almost Sure Convergence Rate
The slowdown factor $r_{\gamma}$ due to the smoothness parameter $\gamma$ appears only with momentum $\beta>0$ and $\gamma<1$. The $r_{\gamma}$ factor is counter-intuitive since momentum is supposed to make convergence faster when the gradient is small. Examining (3.13), the factor $\frac{\beta}{1-\beta}$ appears as a scalar multiplier to the ‘slower’ converging terms, which slows down the overall rate for $\beta>0$.
2.3. Convergence Rate with High Probability
Observe that the inner product term on the right hand side of (2.5) is a martingale difference. If we sum both sides of (2.5) from $t$ to T1T-1, then we get
$F(w_{T})-F_{*}$ $\leq[F(w_{t})-F_{*}]+\sum_{s=t}^{T-1}\alpha_{s}\bigg(\langle\nabla F(w_{s}),\,\delta m_{s}\rangle-\|\nabla F(w_{s})\|^{2}\bigg)$ (2.7)
$+\frac{L}{1+\gamma}\sum_{s=t}^{T-1}\alpha_{s}^{1+\gamma}\,\|\nabla\ell(Z_{s},w_{s})\|^{1+\gamma}.$
The term $M_{T}:=\sum_{s=t}^{T-1}\alpha_{s}\langle\nabla F(w_{s}),\,\delta m_{s}\rangle$ is now a martingale which can be estimated by a term smaller than $\sum_{s=t}^{T-1}\alpha_{s}\|\nabla F(w_{s})\|^{2}$ with high probability using martingale concentration inequalities of Bernstein and Azuma-Hoeffding. These concentration inequalities require strong uniform bound on the martingale difference, which assuming that $\nabla F$ is Lipschitz, can be estimated by Cauchy-Schwarz $|\langle\nabla F(w_{s}),\,\delta m_{s}\rangle|\leq\|w_{s}-w_{*}\|\cdot\|\delta m_{s}\|$. Therefore, such an approach requires estimating $\|w_{s}-w_{*}\|$ and $\|\delta m_{s}\|$ either almost surely or with high probability (lemma 4.1 and corollary 4.7). To produce uniform estimates for these quantities to apply concentration inequalities, we require that the step size be ‘polynomially bounded’, i.e. there exists $c_{1},c_{2}>0$ such that $c_{1}t^{-p}\leq\alpha_{t}\leq c_{2}t^{-p}$ for all $t>0$ (cf. table 2).
There is a nice intuition associated to this approach. Compared to a deterministic algorithm where the error $F(w_{t})-F_{*}$ decreases at every iteration [4, thm 3.4, lemma 2.28], the effectiveness of a stochastic algorithm is shown over several iterations $[t,T]$. That is, one must wait until inefficiencies due to noise ‘averages out’ over several iterations. Therefore, estimating $F(w_{t})-F_{*}$ as a sum over an interval as in (2.7) can be meaningful. We remark that with existing assumptions (2.1 and 2.2), it is difficult to show directly that $F(w_{t})-F_{*}$ decreases almost surely by conventional means (e.g. Borel-Cantelli), which is a key to achieve an optimal convergence rate of $O(t^{-1/2})$ shown in [1] (refer to remark 3.3).